OpenStudy (anonymous):

What are the directrices of the ellipse given by the equation x^2/25 + y^2/9 =1?

OpenStudy (dumbcow):

a^2 = 25 b^2 = 9 For ellipse a^2 = b^2 + c^2 c^2 = 25 - 9 = 16 c = 4 The eccentricity e is c/a e = 4/5 The directrix is 1/e away from edge of ellipse 1/e = 5/4 center is (0,0) edges of major axis at (-5,0) and (5,0) directrices at x = +- 25/4